### 09 Mar 2026 v1

Abstract
We propose a model of whataboutism — a rhetorical strategy that deflects criticism by citing similar misconduct that goes uncriticized on the critic’s side — and study its implications for social norms governing offensive speech. In an infinite-horizon psychological game with two rival camps, agents weigh the intrinsic benefit of offensive speech against the risk of condemnation. External criticism can be deflected via an equilibrium-based whataboutism rebuttal. We characterize the unique dynamically stable Psychological Subgame Perfect Equilibrium and show that the availability of whataboutism exacerbates offensive speech, to the extent that civility norms can break down entirely, especially in polarized societies.
1 Introduction
A fundamental tenet of political liberalism is that even when different factions in society have conflicting ends, they should be able to agree on the legitimate means toward these ends. As Sunstein (2025) writes: “Liberals want people who disagree with each other to find a way to live together, and to smile — or at least to nod respectfully — at their differences.” For example, all sides should agree that public demonstrations are acceptable as long as they do not deteriorate into vandalism. Likewise, making sarcastic comments about a political opponent may be fine, but humiliating her family members should be off-limits — even if one believes that the humiliation is an effective tactic.
In practice, this norm is often enforced in a decentralized manner, via social sanctions that one individual imposes on another. For example, if one politician mocks an opponent’s family members, she will be condemned by politicians or by social-media users. Such condemnations impose a social cost on the offender. They can come from all sides of the political aisle, but intuitively, they carry more weight when they come from the politician’s own camp. Moreover, failure to condemn inappropriate behavior by a member of one’s own camp is itself a potential cause for condemnation. Agents will take these anticipated costs into account when choosing whether to engage in (or tacitly condone) offensive expression.
Whataboutism is a tactic that people sometimes use to deflect a criticism against violations of social norms. Merriam-Webster dictionary defines whataboutism as “the act or practice of responding to an accusation of wrongdoing by claiming that an offense committed by another is similar or worse.” For example, when a prominent public figure from camp A is reprimanded by a member of camp B for mocking the spouse of a camp-B politician (or for keeping silent when another prominent member of camp A did so), a whatboutism response can take the following form: “Why are you criticizing me? Remember when a prominent member of your party ridiculed the child of a politician from my camp? I don’t recall seeing anybody from your party condemning her.”
Whatboutism is a modern-day instance of an ancient rhetorical device known as tu quoque — which Merriam-Webster dictionary defines as “a retort charging an adversary with being or doing what the adversary criticizes in others.” There is a popular sentiment that whatboutism has become increasingly common in recent decades (e.g., see Dykstra (2020)), and that it has sullied public discourse and eroded the liberal ethos that we can agree over means even when we disagree over ends. In Sucio (2025), the Forbes journalist Peter Sucio wrote: [^1] https://www.forbes.com/sites/petersuciu/2025/04/04/whataboutism-has-made-civil-debate-on-social-media-nearly-impossible/ “No matter the topic – especially those about politics and world affairs – it is impossible to have any civil debate on social media without someone responding by making a counter-accusation. This isn’t new, but it has taken such a deep foothold in American politics in recent years, and it isn’t hard to see why.” Sucio went on to attribute the spread of whataboutism to the rise of social media and political polarization.
We propose a formalization of whataboutism as an equilibrium phenomenon, and use this formalism to explore the driving forces behind the phenomenon and its implications for public discourse. In constructing our model, we are guided by three desiderata. First, context should matter: Insulting a soldier on Memorial Day is more offensive than on normal days, and whataboutism arguments would make use of this distinction. Second, the model should enable criticism to be directed not only at offensive speech, but also at subsequent failures to condemn it from within the offender’s camp. Finally, availability of the whatboutism argument should be endogenous: The scenario referred to by one camp’s whataboutism argument is drawn from past speech acts performed by agents on the other camp, which in turn obeyed the same norms shaped by whataboutism. This means that in our model, the social cost of offensive speech should be a function of agents’ equilibrium beliefs regarding the behavior of others.
Accordingly, we construct a model that falls into the general class of extensive-form psychological games (Geanakopolos et al. (1989), Battigalli and Dufwenberg (2009,2022)). Its basic outline is as follows. Society is divided into two symmetric camps. There is a set of states. In each state, an agent from a specific camp is selected to choose whether to perform a speech act that gives him an idiosyncratic intrinsic payoff but offends the other camp. The state captures the context that determines sensitivity to offensive speech. States are ranked by sensitivity: In higher-sensitivity states, agents are less likely to derive much satisfaction from offensive speech and more likely to be insulted by it. [^2] See Gift and Lastra-Anadón (2026) for experimental investigation of how context affects sensitivity to offensive speech.
An agent who considers performing or tacitly supporting an offensive speech act weighs the intrinsic payoff it provides against the cost of being condemned for it. The condemnation can arrive from the rival camp or from one’s own camp. Agents have repeated opportunities to voice disdain for the offensive speech act. When agents fail to condemn one of their own, they are themselves exposed to condemnation in the same way — capturing statements such as: “Shame on you for keeping silent when a member of our/your camp mocked an opponent’s spouse.” The game ends as soon as a condemnation takes place.
A condemnation’s effectiveness depends on whether it is internal or external. Condemnation by a member of one’s own camp is always effective; it automatically inflicts a cost on an agent who performed/supported offensive speech. In contrast, external condemnation can be rebuffed with a whataboutism argument. The argument is based on drawing a play path from the distribution induced by equilibrium strategies, in a state that satisfies three requirements: $(1)$ the state is at least as sensitive as the current one; $(2)$ the initial speech-act choice is faced by an agent from the rival camp; and $(3)$ that agent performs an offensive speech act and meets no internal rebuke along the play path. This definition captures whataboutism arguments, which take the following general form: “In a different context that is at least as sensitive as the present one, someone from your camp performed an offensive speech act and no one from your camp condemned it.” The agent’s cost from external condemnation is defined as the probability that he cannot deflect it with whatboutism. The fact that the cost is determined by the ex-ante equilibrium distribution over play paths is what turns our model into a psychological game.
We characterize the set of Psychological Subgame Perfect Equilibria (PSPE) in this model. In PSPE, players follow a symmetric, stationary strategy, such that whataboutism arguments only invoke a state of equivalent sensitivity. In each state, there is a history-independent propensity to avoid (or internally condemn) offensive speech. This propensity increases with the state’s sensitivity. When sensitivity is below some threshold, the propensity drops to zero: There is a complete breakdown of norms, such that agents always engage in offensive speech in that state, face no internal condemnation, and always use whataboutism to deflect external criticism. A dynamic stability criterion selects a unique PSPE: the one with the lowest amount of offensive speech (i.e., the lowest sensitivity threshold below which norms break down). As society becomes more polarized (in the sense that agents on both camps feel more strongly about expression), offensive speech and the use of whataboutism become more prevalent.
We compare PSPE to a benchmark model in which whataboutism is not allowed. In that model, every state sees a positive fraction of agents who avoid (and internally condemn) offensive speech. These findings are consistent with the intuition that whataboutism erodes the civility of public discourse, and more so in a polarized society.
2 A Model
There are two rival camps, $1$ and $2$, each consisting of a measure one of agents. We use i-i to denote the rival of camp $i$. There is a set of states $S=S_{1}\cup S_{2}$, where $S_{i}=\left\{s_{i}^{1},...,s_{i}^{n}\right\}$ and $S_{1}\cap S_{2}=\emptyset$. Denote $I(s)=i$ when $s\in S_{i}$. In state $s$, camp $I(s)$ gets an opportunity to engage in offensive speech against the other camp. An offensive speech act in state $s$ gives intrinsic utility $v\sim U[0,g_{m}]$ to a member of camp $I(s)$, and intrinsic disutility drawn from $U[0,b_{m}]$ to a member of camp I(s)-I(s). Assume $1<g_{n}<\cdots<g_{1}<2$ and $1<b_{1}<\cdots<b_{n}<2$. A larger $m$ represents greater sensitivity to offensive speech: It generates lower intrinsic utility for offenders and higher disutility for victims. When $m^{\prime}\geq m$, we sometimes write $s_{i}^{m^{\prime}}\succsim s_{j}^{m}$ for any $i,j$, and say that $s_{i}^{m^{\prime}}$ is at least as sensitive as $s_{j}^{m}$.
We now describe an infinite-horizon, sequential-move psychological game in which actions are speech acts. In stage 0, a state $s$ is publicly drawn from the uniform distribution over $S$. In stage $1$, an agent from camp $I(s)$ is randomly selected, privately observes his $v$, and chooses a publicly observed action $a\in\{0,1\}$, where $a=1$ means that the agent performs an offensive speech act (often referred to as “the act”). At every subsequent stage $k>1$, an agent is randomly and uniformly selected from either camp; his valuation is newly and independently drawn. Agents’ action set remains $\{0,1\}$, but now $a=1$ means supporting previous agents, while $a=0$ means condemning them. The game is terminated as soon as an agent chooses $a=0$.
Agents from camp I(s)-I(s) follow an exogenous rule (which we endogenize below): Condemning camp $I(s)$ with probability $\lambda b_{s}$, where $\lambda\in(0,\frac{1}{2})$. The only endogenous objects are the strategies of camp-$I(s)$ agents. Since agents from the rival camp I(s)-I(s) are selected at each stage with probability $\frac{1}{2}$ and terminate the game with probability $\lambda b_{s}$, it follows that any profile $\sigma$ of strategies by camp-$I(s)$ agents induces a well-defined probability distribution over finite terminal histories. For every $s\in S$, let $\alpha_{\sigma}(s)>0$ denote the ex-ante probability that the agent who terminates the game in $s$ is from camp I(s)-I(s), under the strategy profile $\sigma$.
To complete the game’s description, we define the agents’ payoffs. As we will see, preferences are not entirely consequentialistic. Rather, they fit the framework of psychological game theory:
  • When an agent in camp $I(s)$ plays $a=0$ in any stage, his payoff is normalized to 0.
  • When the agent plays $a=1$ and his immediate successor also plays $a=1$, his payoff is $v$. The interpretation is that the agent derives an intrinsic utility from performing/supporting the offensive speech act against the rival camp, and incurs no social cost because he does not face immediate condemnation.
  • When the agent plays $a=1$ and his immediate successor is from his own camp $I(s)$ and plays $a=0$, the agent’s payoff is v1v-1. The interpretation is that in addition to the intrinsic “warm glow” from performing/supporting the offensive speech act, the agent experiences a social sanction in the form of immediate condemnation from his own camp.
  • Finally (this is where the psychological-game component enters), when the agent plays $a=1$ and his immediate successor is from the rival camp I(s)-I(s) and plays $a=0$, the agent’s payoff is defined relative to a fixed strategy profile $\sigma$. Specifically, his payoff is $v$ if he can mount a “rebuttal”, and v1v-1 if he cannot. A rebuttal consists of a choice of a state $s^{\prime}\succsim s$ for which $I(s^{\prime})=-I(s)$, and a terminal history in that state — drawn randomly from the distribution over terminal histories induced by $\sigma$ — such that according to $\sigma,$ the stage-$1$ agent in that history plays $a=1$, and all subsequent agents from camp $I(s^{\prime})$ play $a=1$ in that terminal history.
The game form captures in stylized fashion the following situation. A member from one camp has an opportunity (in a mainstream-media interview, or on social media) to verbally attack the rival camp. This agent may derive intrinsic satisfaction from this attack. Alternatively, the attack may advance the camp’s goal, or energize “the base”. When the agent seizes the opportunity and carries out the verbal attack, there are repeated subsequent opportunities for members from either camp to condemn the attack or tacitly condone it by keeping silent. Like the original attack, these opportunities can involve interviews on traditional media or high-visibility social media posts. The assumption that $g_{m}>1$ for every $m$ means that there is a positive measure of “fanatics”, who are willing to perform/support offensive speech acts against the rival camp, regardless of the state’s sensitivity, even if they face certain condemnation.
The interpretation of the psychological-game component is as follows. When an agent from camp I(s)-I(s) criticizes an agent from camp $I(s)$ for performing/supporting an offensive speech act, the latter agent can employ a “whataboutism” counterargument: In circumstances that are at least as sensitive as $s$, an agent from the critic’s own camp engaged in offensive speech and faced no subsequent internal rebuke. The psychological-game component arises because the rebuttal is based on a random draw from the distribution over play paths induced by the strategy profile $\sigma$. The rebutting agent needs to retrieve a concrete scenario that would arm him with the whatboutism argument. In this scenario, the state needs to be at least as sensitive as $s$; an agent from the rival camp I(s)-I(s) needs to perform an offensive speech act; and in the ensuing play path, there needs to be no internal condemnation, such that the agent who terminates the game is from camp $I(s)$.
Our modeling of the agent’s retrieval process is based on two desiderata: We want to assume that the agent has some control over the retrieval, but also that success is uncertain. Consequently, our model mixes elements of “directed” and “random” search, in analogy to the macro-theory literature (see Wright et al. (2021)). The agent deliberately targets a specific state, but then he — or someone from his own camp who argues on his behalf — randomly retrieves from memory a concrete state-specific scenario. Thus, when the agent plays $a=1$, he does not know for sure whether he will be able to successfully use whatboutism to rebuff out-group condemnation, but his choice of which state to target is optimal in this regard.
The interpretation of $a$ as a speech act
Why do we interpret $a$ as a speech act, rather than as a physical action? Imagine that the whatboutism rebuttal were infeasible. In this case, we would be able to interpret $a=1$ as a physical action that benefits one’s own camp and harms the rival camp. An agent who takes this action incurs a cost if his immediate successor plays $a=0$. This is a fairly standard “network externality”: The agent’s benefit from $a=1$ depends on whether other agents take this action, too.
The whataboutism rebuttal, which enables an agent to save the cost, is hard to interpret except as a verbal argument that has to do with condemnations of bad actions. Thus, the feasibility of whataboutism makes it more natural to interpret actions at stages $k>1$ as verbal acts of condemnation and (explicit or tacit) support. In principle, agents’ choices at stage $k=1$ couldcould be interpreted as physical actions as well as speech acts. Given our assumption that the payoff structure is stationary and does not distinguish between $k=1$ and $k>1$, we find it more natural to restrict attention to the speech-act interpretation of actions at all stages of the game.
“Microfounding” the probability of external criticism
The rule that determines the probability of external condemnation can be endognized as follows. In state $s$, members of camp I(s)-I(s) incur a stochastic cost $c$ when they condemn an offensive speech act performed or supported by members of camp $I(s)$. The cost captures the difficulty of reaching out from one camp to another. Assume that $c\sim U[0,\bar{c}]$, where $\bar{c}>b_{n}.$ Assume that a member of camp I(s)-I(s) condemns the verbal attack from camp $I(s)$ if and only if his disutility from the attack exceeds the cost $c$. Recall that in state $s$, $u\sim[0,b_{s}]$. Therefore, the aggregate probability that camp I(s)-I(s) will condemn offensive speech by camp $I(s)$ is
$\frac{1}{\bar{c}b_{s}}\int_{0}^{b_{s}}\int_{0}^{u}dcdu=\frac{b_{s}}{2\bar{c}}=\lambda b_{s}$
where $\lambda=1/2\bar{c}$. Note that $\lambda<\frac{1}{2}$ since $\bar{c}>b_{n}>1.$
3 A Benchmark
We begin our analysis with a benchmark model in which agents cannot use whataboutism to rebuff criticisms. In this case, an agent who performs an offensive speech act incurs a cost of $1$ whenever his immediate successor (regardless of his affiliation) issues a condemnation. Payoffs in this variant are entirely consequentialistic, and thus do not require the psychological-games machinery. Moreover, in this variant, analysis can be conducted for each state in isolation, whereas the main version of our model creates a linkage across states via the whataboutism counterargument. This comparison will enable us to explore the role of whataboutism in shaping the norms of public discourse.
In what follows we assume that when an agent of camp $I(s)$ is indifferent between $a=0$ and $a=1,$ he chooses $a=1.$
Proposition 1
There is a unique Subgame Perfect Equilibrium in the benchmark model. In each state $s$, each camp-$I(s)$ agent plays $a=1$ if and only if his intrinsic payoff $v$ is above the cutoff
$v_{s}^{\ast}=\frac{\lambda g_{s}b_{s}}{2g_{s}-1}$ (1)
Proof. Since we can analyze equilibrium behavior state by state, we will omit the $s$ subscripts. The proof proceeds in three steps.
Step 1. At each stage, agents follow a cutoff rule.
We show that for every $k>0$, there is a cutoff $v^{k}\in[0,g]$ such that a member of camp $I(s)$ plays $a=1$ if and only if $v\geq v^{k}$. To see why, recall that agents do not observe past agents’ valuations. Stage-$k$ agent’s belief regarding the behavior of stage-$(k+1)$ agent is independent of agent $k$’s valuation. Therefore, if a stage-$k$ agent with valuation $v$ plays $a=1$ ($a=0$), then the same action is also optimal for a agent with valuation $v^{\prime}>v$ ($v^{\prime}<v$). $\square$
Step 2. $v^{k}\in(0,g)$ for every $k$.
Suppose $v^{k}=0$. This means that camp-$I(s)$ agents with $v<\frac{1}{2}\lambda b$ play $a=1$. Such agents earn a negative payoff, regardless of $v^{k+1}$, a contradiction. Now suppose $v^{k}=g$. This means that camp-$I(s)$ agents with $v>1$ play $a=0$, even though playing $a=1$ would generate a strictly positive payoff, a contradiction. $\square$
[…]
$v^{k}=\frac{1}{2}\lambda b+\frac{1}{2}\frac{v^{k+1}}{g}$ (2)
Step 3. The only solution to the system of equations given by (2) is stationary and given by (1).
Rearrange (2) as follows:
$v^{k+1}-v^{k}=(2g-1)v^{k}-\lambda bg$ (3)
Note that 2g1>02g-1>0. If $v^{k+1}-v^{k}>0$ $(<0$) for some $k$, then the sequence $v^{k},v^{k+1},v^{k+2},$... is a divergent increasing (decreasing ) sequence, which is a contradiction because $v$ is bounded between $(0,g)$. It follows that the only solution to (1) is $v^{k+1}=v^{k}$ for every $k$, which yields (1). $\square$
Note that since $\lambda<\frac{1}{2}$, $g>1$ and $b<2$, the R.H.S. of (1) is below $g$, hence $v^{\ast}$ is indeed an interior cutoff given by an indifference condition. $\blacksquare$
The equilibrium fraction of camp-$I(s)$ agents who do not perform/support offensive in state $s$ is
$x_{s}=\frac{\lambda b_{s}}{2g_{s}-1}$ (4)
This fraction is always in $(0,1)$. The fact that $x_{s}>0$ for every $s$ means that norms of public discourse never break down entirely — i.e., there is always a positive fraction of agents who refrain from offensive speech or criticize it internally. Note that $x(s)$ is increasing in the parameter $b_{s}$ that determines the disutility from offensive speech that members of camp I(s)-I(s) incur. It is also increasing in the parameter $\lambda$ that affects camp-$I(s)$ agents’ exposure to external criticism. Finally, it is decreasing in the parameter $g_{s}$ that determines the intrinsic utility that camp-$I(s)$ agents derive from verbally attacking the rival camp.
Note that the R.H.S. of (2) can be interpreted as a standard social-preference effect: An agent trades off the intrinsic utility from bad behavior (offensive speech) and the social cost in the form of frequency with which he is condemned for such behavior. Thus, the benchmark model is essentially a standard model of pro- or anti-social behavior in the presence of social preferences.
4 Analysis: Whataboutism
This section characterizes equilibrium behavior in the main model. As a preliminary step, we derive an expression for agents’ psychological expected payoffs. Fix a strategy profile $\sigma$ and consider a state $s$. Consider a camp-$I(s)$ agent with valuation $v$, who is selected to move in some stage $k$. If the agent chooses $a=0$ and thus terminates the game, his payoff is by definition 0. Suppose now that the agent chooses $a=1$. Let $\beta_{\sigma}$ denote the probability that the stage-$(k+1)$ agent will choose $a=0$, given the history, conditional on being from the same camp $I(s)$. Recall that $\alpha_{\sigma}(s^{\prime})$ is the ex-ante probability (induced by $\sigma$) that the agent who terminates the game in $s^{\prime}$ is from camp $I(s^{\prime})$. This is the probability that a randomly sampled state-$s^{\prime}$ play path will support a whataboutism rebuttal by a member of camp $I(s^{\prime})$. Then, the stage-$k$ agent’s psychological expected payoff given $\sigma$ from playing $a=1$ is
$v-\frac{1}{2}\beta_{\sigma}-\frac{1}{2}\lambda b_{s}\cdot\min_{s^{\prime}\succsim s\mid I(s^{\prime})=-I(s)}(1-\alpha_{\sigma}(s^{\prime}))$ (5)
The first term in this expression is the agent’s intrinsic payoff from performing/supporting the speech act. The second term represents the probability that the agent faces condemnation from within his own camp, which by assumption is impossible to deflect. The third term represents the probability that the agent faces external condemnation which he cannot rebut with a whatboutism argument. Note that the minimum operator is due to the agent’s ability to select among the eligible states (i.e., those that are at least as sensitive as $s$) to maximize the probability of being able to say: “But what about state $s^{\prime}$? An agent from your camp performed an offensive speech act in that state, which is at least as sensitive as $s$, and I didn’t see anyone from your camp criticizing him.”
Note that the final term is calculated with respect to the strategy profile $\sigma$. As a result, the agent’s expected payoff is defined in terms of the ex-ante strategy profile, rather than the continuation strategy given the history at which the agent moves. This is a hallmark of psychological-game expected payoff calculations.
We now introduce our notion of equilibrium, due to Geanakopolos et al. (1989).
Definition 1
A strategy profile $\sigma$ is a Psychological Subgame Perfect Equilibrium (PSPE) if at every history, each agent chooses $a=0$ if and only if it maximizes his psychological expected payoff given $\sigma$.
The sequential rationality requirement is exactly as in conventional SPE; the only difference is the dependence of agents’ expected payoff from $a=1$ on the ex-ante strategy profile, which includes behavior in other states.
A two-state example
Let $S=\{s_{1},s_{2}\}$ where $s_{1}\sim s_{2}$, such that both states are associated with the same $(g,b)$. We focus on stationary and symmetric PSPE — i.e., there is a cutoff $v^{\ast}$ such that at every $s$, a camp-$I(s)$ agent with valuation $v$ chooses $a=1$ at any history, if and only if $v\geq v^{\ast}$. In such an equilibrium, $\beta$ (the probability that a camp-$I(s)$ agent plays $a=0$) is given by $v^{\ast}/g$, independently of the history. Therefore,
$\alpha_{\sigma}(s)=\left(1-\frac{v^{\ast}}{g}\right)\cdot\frac{\frac{1}{2}\lambda b}{\frac{1}{2}\lambda b+\frac{1}{2}\frac{v^{\ast}}{g}}$ (6)
The second term is the conditional probability that given a terminal history, the last move is by an agent from camp I(s)-I(s). This probability is calculated according to each camp’s stationary probability of playing $a=0$. Note that the expression for $\alpha_{\sigma}(s)$ is non-linear in the cutoff $v^{\ast}$ — unlike the benchmark case. This non-linearity reflects the fact that the whataboutism argument is based on the distribution over play paths induced by agents’ ex-ante strategies, and it will play an important role in the sequel.
Rearranging (6), we obtain
$1-\alpha_{\sigma}(s)=\frac{v^{\ast}+b\lambda v^{\ast}}{v^{\ast}+b\lambda g}$
The cutoff $v^{\ast}$ equates the agent’s payoff from $a=1$ and $a=0$:
$v^{\ast}-\frac{1}{2}\frac{v^{\ast}}{g}-\frac{1}{2}\lambda b\frac{v^{\ast}+b\lambda v^{\ast}}{v^{\ast}+b\lambda g}=0$
This equation has potentially two solutions. One solution is $v^{\ast}=0$, which corresponds to a stationary equilibrium in which all agents always play $a=1$ regardless of their valuation. When $g<1+\frac{1}{2}b\lambda$, there is an additional solution:
$v^{\ast}=\frac{\lambda bg}{2g-1}\left(\lambda b-2g+2\right)$ (7)
The fraction of camp-$I(s)$ agents who play $a=0$ in state $s$ at any stage is $x^{\ast}=v^{\ast}/g$. Note that since $g>1$ and $b\lambda<1$, this fraction $x^{\ast}$ is equal to the benchmark level (4) multiplied by a “discount factor” $b\lambda-2g+2<1$. This factor captures an equilibrium effect of whataboutism, which raises the fraction of agents who perform/support offensive speech acts. When $b\lambda-2g+2<0$ there is no interior solution, and $x^{\ast}=0$ in the unique equilibrium — i.e., all norms that constrain public expression break down.
4.1 The Main Result
In this subsection we characterize the PSPE and show that there is a unique stable equilibrium. The following notation will be useful:
$c_{m}$ $=$ $\frac{\lambda b_{m}}{2g_{m}-1}$ (8)
$\theta_{m}$ $=$ $\lambda b_{m}-2g_{m}+2$
$\phi_{m}(z)$ $=$ $\frac{\lambda b_{m}z+z}{\lambda b_{m}+z}$
Note that $c_{m}$ is equal to the fraction of agents who play $a=0$ in SPE under the benchmark model. Note also that $\theta_{m}<1$ (since $\lambda b_{m}<1$ and $g_{m}>1$) and increases in $m$. Let $M$ be the lowest value of $m$ for which $\theta_{m}\geq 0$.
Proposition 2
For every $m^{\ast}\in\{M,...,n+1\}$, there is a PSPE such that: (i) for every $m<m^{\ast}$, camp-$I(s_{i}^{m})$ agents always play $a=1$ in state $s_{i}^{m}$; (ii) for every $m\geq m^{\ast}$, a camp-$I(s_{i}^{m})$ agent with valuation $v$ plays $a=1$ if and only if $v\geq v_{m}^{\ast}=c_{m}\theta_{m}g_{m}$. There are no other equilibria.
Proof. The proof proceeds stepwise.
Step 1: In PSPE, agents follow a stationary cutoff strategy in each state.
By essentially the same arguments as in Step 1 of the proof of Proposition 1, for very state $s$ and every stage $k$, there is a cutoff $v_{s}^{k}$ such that a camp-$I(s)$ agent with valuation $v$ who moves in stage $k$ plays $a=1$ if and only if $v\geq v_{s}^{k}$.
Given a PSPE $\sigma,$ $\mu_{s}\left(\sigma\right):=\min_{s^{\prime}\succsim s\mid I(s^{\prime})=-I(s)}(1-\alpha_{\sigma}(s^{\prime}))$ is the probability that external condemnation of camp $I(s)$ in state $s$ cannot be rebuffed by whatboutism. Then, in state $s$, a camp-$I(s)$ agent with valuation $v$ who moves in stage $k$ earns a payoff of
$v-\frac{1}{2}\lambda b_{s}\mu_{s}\left(\sigma\right)-\frac{1}{2}\frac{v_{s}^{k+1}}{g_{s}}$ (9)
from $a=1$, and a payoff of zero from $a=0$. If $\mu_{s}\left(\sigma\right)>0$, then by the same arguments as in steps 2 and 3 of the proof of Proposition 1, $v_{s}^{k}=v_{s}^{k+1}$ for every $k$. If $\mu_{\sigma}(s)=0$, then by the definition of the cutoff, $v_{s}^{k+1}=2g_{s}v_{s}^{k}$, implying that the sequence $v_{s}^{k},v_{s}^{k+1},$... is a divergent increasing sequence. But this violates the assumption that $v\leq g_{s}$. $\square$
Let $v_{s}^{\ast}$ denote the cutoff that characterizes state $s$. Fix $i=1,2$. When $s=s_{i}^{m}$ ($s=s_{-i}^{m}$), we use $x_{m}=v_{s}^{\ast}/g_{s}$ ($y_{m}=v_{s}^{\ast}/g_{s}$) to denote the fraction of members of camp $i$ (i-i) who play $a=0$ in equilibrium in state $s$.
Step 2: In every state $s_{i}^{m}$,
$\mu_{s_{i}^{m}}\left(\sigma\right)=\min_{m^{\prime}\geq m}\phi_{m^{\prime}}(y_{m^{\prime}})$ (10)
By Step 1, in equilibrium, every state $s$ is associated with a stationary cutoff $v_{s}^{\ast}$ that governs camp-$I(s)$ agents’ behavior. Therefore, for every state $s^{\prime}$ where $I(s^{\prime})\neq I(s)$, we obtain essentially the same expression as (6), repeated here for convenience:
$\alpha_{\sigma}(s^{\prime})=\left(1-\frac{v_{s^{\prime}}^{\ast}}{g_{s^{\prime}}}\right)\cdot\frac{\frac{1}{2}\lambda b_{s^{\prime}}}{\frac{1}{2}\lambda b_{s^{\prime}}+\frac{1}{2}\frac{v_{s^{\prime}}^{\ast}}{g_{s^{\prime}}}}$
It follows that for $s^{\prime}=s_{-i}^{m^{\prime}},$ the probability $1-\alpha_{\sigma}(s^{\prime})$ simplifies to the expression $\left(\lambda b_{m^{\prime}}y_{m^{\prime}}+y_{m^{\prime}}\right)/\left(\lambda b_{m^{\prime}}+y_{m^{\prime}}\right),$ which by the definition of $\mu_{s}\left(\sigma\right)$ yields equation (10). $\square$
Step 3: PSPE can be characterized as follows: For every m=1,,nm=1,...,n,
$x_{m}$ $=$ $c_{m}\min_{m^{\prime}\geq m}\phi_{m^{\prime}}(y_{m^{\prime}})$ (11)
$y_{m}$ $=$ $c_{m}\min_{m^{\prime}\geq m}\phi_{m^{\prime}}(x_{m^{\prime}})$
Consider a state $s=s_{i}^{m}.$ From Step 1 and equation (9) it follows that the equilibrium cutoff at this state satisfies $v_{s}^{\ast}=\frac{1}{2}\lambda b_{s}\mu_{s}\left(\sigma\right)+\frac{1}{2}(v_{s}^{\ast}/g_{s})$. Dividing both sides of this equation by $g_{s}$, replacing $(b_{s},g_{s})$ with $(b_{m},g_{m})$ and rearranging yields
$x_{m}=\frac{\lambda b_{m}}{2g_{m}-1}\mu_{s_{i}^{m}}\left(\sigma\right)$
which coincides with the stated equation for $x_{m}$ by the definitions of $c_{m}$ and Step 3. An analogous computation for a state $s_{-i}^{m}$ yields the desired equation for $y_{m}.$ $\square$
The final step of the proof uses the following lemmas.
Lemma 1
For every m=1,,n,m=1,...,n, and every $z>0$: $(i)$ $\phi_{m}(c_{m}z)>z$ if $z<\theta_{m}$, and $\phi_{m}(c_{m}z)<z$ if $z>\theta_{m}$; and $(ii)$ $\phi_{m}(c_{m}z)=z$ has the unique solution $z=\theta_{m}.$
Proof. Note that
$\phi_{m}(c_{m}z)-z=\frac{\left(1+\lambda b_{m}\right)c_{m}z}{\lambda b_{m}+c_{m}z}-z=\frac{c_{m}z\left[\left(\lambda b_{m}-2g_{m}+2\right)-z\right]}{\lambda b_{m}+c_{m}z}$
Since $\lambda b_{m}+c_{m}z>0$ and $c_{m}z>0$, the result follows. $\square$
Lemma 2
Suppose there exist $m\in\left\{1,...,n\right\}$ and $T\geq\theta_{m}$ such that
$x_{m}$ $=$ $c_{m}\min\left\{\phi_{m}(y_{m}),T\right\}$ (12)
$y_{m}$ $=$ $c_{m}\min\left\{\phi_{m}(x_{m}),T\right\}$
If $\theta_{m}>0$, then either $x_{m}=y_{m}=0$ or $x_{m}=y_{m}=c_{m}\theta_{m}$ (and in the latter case, $\phi_{m}(x_{m})=\phi_{m}(y_{m})=\theta_{m}$). If $\theta_{m}\leq 0$, then $x_{m}=y_{m}=0$.
Proof. Let $(x_{m},y_{m})$ be the solutions to (12). Since $\phi_{m}\left(\cdot\right)$ is strictly increasing on $0,\infty),$ $x_{m}>y_{m}$ implies $\phi_{m}\left(x_{m}\right)>\phi_{m}(y_{m})$ and hence, $\min\left\{\phi_{m}(x_{m}),T\right\}\geq\min\left\{\phi_{m}(y_{m}),T\right\}$, which implies $y_{m}\geq x_{m},$ a contradiction. Similarly, $y_{m}>x_{m}$ is impossible. Therefore, $x_{m}=y_{m}=t$, such that $t=c_{m}\min\left\{\phi_{m}(t),T\right\}$. Note that since $\phi_{m}(0)=0$, $t=0$ solves this equation.
Let us now look for strictly positive solutions for $t$. Suppose $\phi_{m}(t)>T$. Then, $t=c_{m}T$, such that $\phi_{m}(c_{m}T)>T.$ By Lemma 1, this holds if and only if $T<\theta_{m}$, a contradiction. Therefore, $\phi_{m}(t)\leq T$, such that $t=c_{m}\phi_{m}(t)$. For $t>0$, this implies $1=c_{m}\left(\lambda b_{m}+1\right)/(\lambda b_{m}+t)$, such that $t=c_{m}\theta_{m}$, and therefore $\phi_{m}(t)=\theta_{m}$. By assumption, $\theta_{m}\leq T$, hence the solution is consistent with the guess that $\phi_{m}(t)\leq T$. We conclude that a strictly positive solution to (12) exists iff $\theta_{m}>0,$ and when it exists it is equal to $c_{m}\theta_{m}.$ $\square$
Step 4: In PSPE, there is $m^{\ast}\in\{M,...,n+1\}$, such that $x_{m}=y_{m}=0$ for $m<m^{\ast},$ while $x_{m}=y_{m}=c_{m}\theta_{m}$ and $\phi_{m}(x_{m})=\phi_{m}(y_{m})=\theta_{m}$ for $m\geq m^{\ast}$.
First, suppose that in some PSPE, $x_{m}=0$ or $y_{m}=0$ for some $m$. Then, since $\phi_{m^{\prime}}(0)=0$ for every $m^{\prime}$, Step 3 implies that $x_{m^{\prime}}=y_{m^{\prime}}=0$ for every $m^{\prime}\leq m$. Thus, in PSPE, there is $m^{\ast}\in\{1,...,n+1\}$ such that $x_{m}=y_{m}=0$ for every $m<m^{\ast}$, and $x_{m},y_{m}>0$ for every $m\geq m^{\ast}$. If $m^{\ast}=n+1$, we are done. Now suppose $m^{\ast}\leq n$. The proof proceeds by backward induction on $m.$
Base case: $m=n$. Clearly, $\min_{m\geq n}\phi_{m}(y_{m})=\phi_{n}(y_{n})$. Apply Lemma 2 with $T=+\infty$. If $\theta_{n}\leq 0$, we have $x_{n}=y_{n}=0$, a contradiction. Therefore, from now on we must assume $\theta_{n}>0$. In this case, the lemma implies $x_{n}=y_{n}=c_{n}\theta_{n}$ and $\phi_{n}(x_{n})=\phi_{n}(y_{n})=\theta_{n}$.
Inductive step. Fix some $m\in\{m^{\ast},...,n\}$, and assume the claim holds for m+1,,n+1m+1,...,n+1. By the inductive step, $x_{m^{\prime}}=y_{m^{\prime}}=c_{m^{\prime}}\theta_{m^{\prime}}$ and $\phi_{m^{\prime}}(x_{m^{\prime}})=\phi_{m^{\prime}}(y_{m^{\prime}})=\theta_{m^{\prime}}$ for al $m^{\prime}>m$. Note that $\theta_{m}$ is increasing in $m$. Therefore,
$\min_{m^{\prime}>m}\phi_{m^{\prime}}(x_{m^{\prime}})=\min_{m^{\prime}>m}\phi_{m^{\prime}}(y_{m^{\prime}})=\theta_{m+1}$
Applying Lemma 12 with $T=\theta_{m+1}\geq\theta_{m}>0$, we obtain $x_{m}=y_{m}=c_{m}\theta_{m}$ and $\phi_{m}(x_{m})=\phi_{m}(y_{m})=\theta_{m}$. This completes the inductive proof. $\square$
It is now straightforward to confirm that the sequences $(x_{m})$ and $(y_{m})$ described by Step 4 solve (11). Therefore, Step 4 fully characterizes the set of PSPE. $\blacksquare$
Proposition 2 means that any PSPE is characterized by a “sensitivity threshold” $m^{\ast}\geq M$. In state $s_{i}^{m},$ $m\geq m^{\ast}$, the probability that a camp-$i$ agent refrains from offensive speech (and condemns his comrades when they do not) is $x_{m}=v_{m}^{\ast}/g_{m}$, which is equal to the corresponding probability in the benchmark ($c_{m}$) multiplied by the “discount factor” $\theta_{m}.$ Since both $c_{m}$ and $\theta_{m}$ increase with $m,$ the probability of condemnation increases with the sensitivity of the state. The term $\theta_{m}$ can be interpreted as the “whataboutism effect”, which lowers the likelihood that agents condemn members of their camp for performing/supporting offensive speech. The whataboutism effect can be dramatic and lead to a “breakdown of norms” in the sense that in if $\theta_{m}<0$ for some state $s_{i}^{m},$ then in all weakly less sensitive states, all agents (regardless of their camp) perform/support offensive speech.
A key aspect of the result is that when an agent employs the whataboutism rebuttal in some state $s_{i}^{m}$, he chooses to target the equally sensitive mirror state $s_{-i}^{m}$. This feature, combined with the symmetries of the model, leads to the symmetry of equilibrium behavior across camps, and to the feature that the equilibrium equations treat each sensitivity level $m$ in isolation. The logic behind this feature is that $\alpha_{\sigma}(s)$ (which is the probability that a randomly sampled state-$s$ play path supports a whataboutism argument by a member of camp I(s)-I(s)) is decreasing in the sensitivity of $s$. Establishing this monotonicity takes the bulk of the proof of Proposition 2. Therefore, a agent who employs whataboutism in state $s_{i}^{m}$ will target at the least sensitive eligible state, which is therefore $s_{-i}^{m}$.
The PSPE are ranked by their amount of offensive speech: The higher $m^{\ast}$, the higher that amount. In particular, there is always an equilibrium with complete breakdown of norms, where $m^{\ast}=n+1$ such that agents always engage in (and never condemn) offensive speech. To see why, note that the equilibrium probability that in state $s$ a camp-$I(s)$ agent chooses $a=0$ is equal to the benchmark probability $\lambda b_{s}g_{s}/(2g_{s}-1)$, multiplied by the minimal probability that this agent will not be able to counter a criticism from the rival camp; and if all agents always choose $a=1,$ this probability is zero. However, as we show below, the unique dynamically stable equilibrium among all PSPE is the one with the least amount of offensive speech.
The proof of Proposition 2 establishes that a PSPE induces a sequence $x=(x_{m})_{m=1}^{n}$ — where $x_{m}$ is the probability that a camp-$i$ agent chooses $a=0$ in state $s_{i}^{m}$ — that satisfies $x_{m}=c_{m}\phi_{m}(x_{m})$ for every m=1,,nm=1,...,n, where $c_{m}$ and $\phi_{m}$ are given by (8). Note that if there is a strictly positive solution to the equation for $m$, then it also solves the equation $x_{m}=\phi_{m}(c_{m}x_{m}).$ This follows from the observation that when $x_{m}>0$, both equations reduce to being linear in $x_{m}.$ We therefore have the following corollary.
Corollary 1
A sequence of probabilities $(x_{m})_{m=1}^{n}$ is consistent with PSPE if for each $m,$ $x_{m}$ solves $x_{m}=\phi_{m}(c_{m}x_{m}).$
This corollary means that we can regard the R.H.S. of the equation $x_{m}=\phi_{m}(c_{m}x_{m})$ as a self-mapping from $x$ to itself, analogous to a best-reply correspondence in standard game theory. The $x_{m}$ on the R.H.S. represents the fraction of agents on either camp who play $a=0$ in the states $s_{1}^{m}$ and $s_{2}^{m}$. The $x_{m}$ on the L.H.S. is induced by the individual agent’s cutoff $v_{m}^{\ast}$, which is his best-response to the population-level behavior. In light of this interpretation of the equilibrium equations, we propose the following notion of dynamic stability.
Definition 2
A PSPE is dynamically stable if its induced sequence $x=(x_{m})_{m=1}^{n}$ satisfies the following: There is $\delta>0$ such that for every $m$ and $x_{m}^{\prime}$ satisfying $\left|x_{m}^{\prime}-x_{m}\right|<\delta$, we have $\left|\phi_{m}(c_{m}x_{m}^{\prime})-\phi_{m}(c_{m}x_{m})\right|<\left|x_{m}^{\prime}-x_{m}\right|$.
Put differently, a PSPE is dynamically stable if at every state, a small perturbation in the fraction of agents (in both camps) who choose $a=0$ in any state leads to a smaller change in the probability that an individual agent chooses that action. [^3] Note that $\phi\left(cx\right)$ is monotone in $x$, so when $x$ changes, $\phi\left(cx\right)$ changes in the same direction.
Proposition 3
The unique dynamically stable PSPE has $m^{\ast}=M$.
Proof. First, let us establish that the PSPE with $m^{\ast}=M$ is dynamically stable. By Lemma 1, for every $m\geq M,$ if there exists a solution $x_{m}>0$ to the equation $\phi_{m}(c_{m}x_{m})=x_{m},$ then it is unique. Moreover, $x_{m}^{\prime}>\phi_{m}(c_{m}x_{m}^{\prime})$ for every $x_{m}^{\prime}>x_{m}$ while $x_{m}^{\prime}<\phi_{m}(c_{m}x_{m}^{\prime})$ for every $x_{m}^{\prime}<x_{m}.$ Hence, for any such solution $x_{m}$,
$\phi_{m}(c_{m}x_{m}^{\prime})-\phi_{m}(c_{m}x_{m})=\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}<x_{m}^{\prime}-x_{m}$
for any $x_{m}^{\prime}>x_{m}$, while
$\phi_{m}(c_{m}x_{m}^{\prime})-\phi_{m}(c_{m}x_{m})=\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}>x_{m}^{\prime}-x_{m}$
for any $x_{m}^{\prime}<x_{m}$. Now consider $m<M$. By the definition of $M$ (as the lowest value of $m^{\prime}$ for which $\theta_{m^{\prime}}>0$), the only non-negative solution to $\phi_{m}(c_{m}x_{m})=x_{m}$ is $x_{m}=\phi_{m}(c_{m}x_{m})=0$; and by Lemma 1, $\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}^{\prime}<0$ for all $x_{m}^{\prime}>0$. Therefore,
$\phi_{m}(c_{m}x_{m}^{\prime})-\phi_{m}(c_{m}x_{m})=\phi_{m}(c_{m}x_{m}^{\prime})<x_{m}^{\prime}=x_{m}^{\prime}-x_{m}$
We have thus established the dynamic stability of the PSPE with $m^{\ast}=M$.
Finally, consider a PSPE with $m^{\ast}>M$. Thus, $x_{M}=0$. By the definition of $M$ and Lemma 1, we can find $x_{m}^{\prime}>0$ such that $\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}^{\prime}>0$, which means that the equilibrium is not dynamically stable. $\blacksquare$
Thus, dynamic stability uniquely selects the PSPE with the lowest amount of offensive speech.
Non-linearity and the uniform distribution of $v$
From a narrowly technical point of view, what makes the equilibrium analysis of whataboutism distinct from the benchmark model, is that the cost of playing $a=1$ is non-linear in the behavior of others. This non-linearity arises from the way the success rate of a whatboutism rebuttal depends on the equilibrium state-dependent frequency of internal condemnation.
This technical comment puts our assumption that $v$ (the intrinsic payoff from offensive speech acts) is uniformly distributed in perspective. This assumption is made partly for tractability, but also to make it clear that the non-linear effects of whataboutism do not rely on distributional assumptions. Our results would carry over to non-uniform distributions of $v$, albeit without transparent closed-form expressions for equilibrium cutoffs.
4.2 The Frequency of Offensive Speech and Whataboutism
We use our characterization of the unique dynamically stable PSPE to analyze the equilibrium frequency of offensive speech and the use of whataboutism. By Proposition 2, the fraction of agents who refrain from offensive speech (or condemn it internally) in the states of sensitivity $m$ is
$x_{m}$ $=$ $y_{m}=\frac{v_{m}^{\ast}}{g_{m}}$
$=$ $\max\left\{0,c_{m}\theta_{m}\right\}$
$=$ $\max\left\{0,\frac{\lambda b_{m}}{2g_{m}-1}\left(\lambda b_{m}-2g_{m}+2\right)\right\}$
The interpretation of the condition $\theta_{m}=\lambda b_{m}-2g_{m}+2>0$ (which holds for $m\geq M$) is that the probability of external condemnation ($\lambda b_{m}$) is more than twice the fraction of “fanatics” in the camp choosing whether to engage in offensive speech ($g_{m}-1$). The following result applies comparative statics to the expression (4.2).
Remark 1
The equilibrium fraction of agents who choose $a=1$ in states of sensitivity $m$ weakly increases in $g_{m}$ and weakly decreases in $\lambda$ and $b_{m}$ (strictly so when $m\geq M$).
The property that $x_{m}$ increases in $b_{m}$ and decreases in $g_{m}$ is shared by the benchmark model. However, the availability of whataboutism introduces a “multiplier effect” given by $\theta_{m}$. The same increase in $g_{m}$, say, leads to a stronger decrease in $x_{m}$, relative to the benchmark. The intuition is as follows. When $g_{m}$ goes up, the fraction of “fanatics” in each camp grows. This has a direct positive effect on the propensity to play $a=1$, as in the benchmark model, because playing $a=1$ exposes an agent to weaker internal criticism. However, unlike the benchmark model, there is also an indirect effect: The increase in the fraction of fanatics in the rival camp also makes it easier to deflect external criticism with whataboutism (because it is easier to come with the scenario that the rival camp engaged in offensive speech and faced no internal rebuke). The combined effect is stronger than in the benchmark model.
Let us now turn to the equilibrium usage of whataboutism. The probability that the whataboutism rebuttal occurs in state $s_{i}^{m}$ is the joint probability of the following two events: $(1)$ the first-moving agent (who belongs to camp $i$) chooses $a=1$, and the agent who terminates the game by condemning him is from camp i-i; and $(2)$ in the equally sensitive state $s_{-i}^{m},$ the first-moving agent (who belongs to camp i-i) chooses $a=1$, and the agent who terminates the game by condemning him is from camp $i$. By symmetry, the probabilities of these two events are equal and given by:
$\left(1-x_{m}\right)\left(\frac{\frac{1}{2}\lambda b_{m}}{\frac{1}{2}\lambda b_{m}+\frac{1}{2}x_{m}}\right)=2g_{m}-\lambda b_{m}-1=1-\theta_{m}$
where the left-hand equality is obtained by plugging (4.2). This implies the following observation.
Remark 2
The frequency of whataboutism in states of sensitivity $m$ is $\min\{1,\left(1-\theta_{m}\right)^{2}\}$. It weakly increases in $g_{m}$ and weakly decreases in $\lambda$ and $b_{m}$ (strictly so when $m\geq M$).
Thus, as sensitivity becomes weaker, we observe more offensive speech and more whataboutism in equilibrium. When norms break down entirely such that all agents engage in offensive speech, whataboutism is always employed.
The Equilibrium Effect of Polarization
Modify the parameters $b$ and $g$ as follows: For every $m$, replace $b_{m}$ and $g_{m}$ with $kb_{m}$ and $kg_{m}$, respectively, where $k\geq 1$. This captures a society that is more polarized in its attitudes to offensive speech: The people who derive satisfaction from offensive speech feel more strongly about it, while their targets are more likely to take offense. Note also that the difference between the two camps expected valuations is $k(g_{m}+b_{m})/2$ (recall that $b_{m}$ indicates a disutility). Hence, polarization in this usual sense also increases with $k$.
Differentiate the equilibrium expressions for $x_{m}$ and the frequency of whatboutism in state $m$ w.r.t $k$, and evaluate the derivative at $k=1$. The derivative of $x_{m}$ is negative while the derivative of $(1-\theta_{m})^{2}$ is positive. This means that as society becomes more polarized, both offensive speech and whatboutism become more frequent.
5 Conclusion
This paper has formalized the argument that whataboutism erodes civility norms in public discourse, and that both offensive speech and the use of whataboutism proliferate in polarized societies . We have argued that psychological game theory offers a suitable tool for modeling whataboutism, because the effectiveness of this rhetorical device depends on the relative equilibrium frequencies of offensive speech acts and the reactions to it on both sides of the political aisle. The psychological-game factor introduces a non-linear dependence on equilibrium strategies into agents’ expected payoff calculations, which generates potential equilibrium multiplicity. However, dynamic stability considerations select a unique equilibrium, which nonetheless is qualitatively distinct from equilibrium in a benchmark model in which anti-social behavior is incentivized by a standard social cost that is linear in the behavior of others.
Although our discussion of the equilibrium effects of whataboutism has suggested that these effects are adverse, we have refrained from explicit utilitarian welfare analysis. Since utilitarianism favors agents who derive sadistic pleasure from offending others or are quick to take offense from controversial speech, we believe it is inappropriate in the present context.